OpenStudy (anonymous):

Graph the inequality. 4x – 2y ≤ -12

OpenStudy (callisto):

1. put x=0 into the equation 4x-2y=12, solve y -2y = 12 => y = -6 2. put y=0 into the equation 4x-2y=12, solve x 4x = 12 => x=3 3. Plot (0, -6) and (3,0) on a graph. Draw a line joining the 2 pointss. The line is NOT dotted line as the inequality sign here is greater than or equal to 4. Put (0, 0)as a test point LHS = 4(0) - 2(0) = 0<12 So, (0, 0) is a point lies on the region. Shade the region where (0, 0) is located

OpenStudy (anonymous):

if you have a graphing calculator...solve for y -2y less than or equal to -4x-12 y is greater than or equal to x+6 hopefully this helps :)